OpenStudy (anonymous):

Using de Moivre’s theorem, express cos 6θ and sin 6θ in powers of cos θ and sin θ. Do i have to use binomial expansion??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i want full explanation

OpenStudy (anonymous):

\[=\cos^{6}+6i \cos^{5}\sin-15\cos^{4}\sin^{2}-20i \cos^{3}\sin^{3}+15\cos^{2}\sin^{4}+6i \cos \sin^{5}-\sin^{6}\].